OpenStudy (anonymous):

Factor completely x 3 +2x 2 -15x A. x (x -3)(x +5) B. -3x (x +5)(x +1) C. x (x +5)(x +3) D. 5x (x+1)(x -3) i need help

OpenStudy (zale101):

Step one: Factor out the common factor of each term \(x^3+2x^2-15x\)

OpenStudy (anonymous):

oh ok i got it now it s b

OpenStudy (zale101):

Do you know how to factor? :)

OpenStudy (anonymous):

ooh

OpenStudy (zale101):

I see x can be factored out in \(x^3+2x^2-15x\) each term consists of x

OpenStudy (anonymous):

yes zale101

OpenStudy (zale101):

so, \(x(x^2+2x-15)\)

OpenStudy (zale101):

we are not done with factoring, you need to factor out the \(x^2+2x-15\)

OpenStudy (anonymous):

its between a and c by what im understanding

OpenStudy (zale101):

\(x^2+2x-15\) \(x^2\)=x*x so, (x )(x ) Nx + Ax = +2x N * A = -15 what two numbers (N and A) that when added you get +2 and when multiplied you get -15 ?

OpenStudy (anonymous):

-3 and positive 5

OpenStudy (zale101):

Yes :)

OpenStudy (anonymous):

oh now i got it thank you for your help zale101

OpenStudy (zale101):

Ok, no problem :)

OpenStudy (zale101):

\(x^3+2x^2+15x\) \(x(x^2+2x-15)\) x(x-3)(x+5)

OpenStudy (anonymous):

sure did help by you helping me out i did get the right answer as you